Leadol) nitrate is added slowly to a solution that is 0.0500 M in Ci ions. Calculate the concentration of Pb ions (in required to initiate the precipitation of Pbci2 molli) (K for PbCI, is 2.40 x 10 4) sp
Answer #1

Pbcl2 = Pb 2+ + 2cl-

[cl-] = 0.05M

[Pb2+] = ksp/[cl-] ^2= 2.4*10-4 / 0.05^2 = 9.6 * 10-2 M

Concentration of Pb 2+ should be greater than 9.6 * 10-2 M to initiate the precipitate